OpenStudy (annah.bee):

Francine and Cheryl received equal scores on a test made up of multiple choice questions and an essay. Francine got 34 multiple choice questions correct and received 14 points for her essay. Cheryl got 30 multiple choice questions correct and received 22 points for her essay. How many points was each multiple choice question worth? _____ points

OpenStudy (danjs):

The Total points earned on the test is then, Total Points = (# multiple choice right)*(Points per multiple choice) + (#essay right)*(Points per essay)

OpenStudy (danjs):

m = points per multipl choice, e = essay points per question Francine - -- Total Earned = 34*(m) + 1*14 Cheryl - - Total Earned = 30*(m) + 1*22

OpenStudy (danjs):

It says they both received the same total points, so set the expressions for both equal and solve for m

OpenStudy (annah.bee):

thanks!

OpenStudy (danjs):

welcome